NASA International Space Station On-Orbit Status 21 September 2015

©NASA

NASA International Space Station On-Orbit Status 21 September 2015.

Japan's resupply ship, the "Kounotori" HII Transfer Vehicle-5 (HTV-5), will end its stay at the International Space Station next Monday. A new Russian Progress cargo mission will launch to replenish the crew Oct. 1. While the HTV-5 is winding down its mission, the six-member Expedition 45 is conducting human research and lab maintenance.

First time space-flyers Kjell Lindgren from NASA and Yui Kimiya Yui from Japan checked each other's eyes today for the Ocular Health study. Lindgren then moved on to a six-month inspection of the treadmill inside the Unity module while Kimiya replaced bacteria filters in the Destiny lab module.

Station Commander Scott Kelly started his workday swabbing his body to collect microorganism samples for the Microbiome study. Kelly later cleaned a science freezer before measuring the acoustic levels inside the space station.

There are three veteran cosmonauts onboard the station with a combined 8 missions, including the current mission, on the orbital lab. Flight Engineer Oleg Kononenko worked in the Microgravity Science Glovebox for the OASIS liquid crystal experiment. Flight Engineer Sergey Volkov with One-Year crew member Mikhail Kornienko studied how living in space affects a crew member's cardiovascular and respiratory systems for the DAN experiment.

On-Orbit Status Report

Ocular Health Testing: Lindgren and Yui have initiated their Flight Day 60 Ocular Health testing by performing vision and tonometry tests, taking blood pressure measurements, and completing a vision questionnaire.

Microbiome Investigation: Kelly completed survey questions and performed body sample collections for his Flight Day 180 session of the Microbiome investigation. Microbiome investigates the impact of space travel on both the human immune system and an individual's microbiome (the collection of microbes that live in and on the human body at any given time).

Observation Analysis of Smectic Islands in Space (OASIS) Sample Exchange: Kononenko performed an OASIS sample exchange today, installing the third of four sample cartridges into the experiment hardware housed in the Microgravity Science Glovebox (MSG). OASIS studies the unique behavior of liquid crystals in microgravity, including their overall motion and the merging of crystal layers known as smectic islands. Liquid crystals are used for display screens in televisions and clocks, and they also occur in soaps and in cell membranes. The experiment allows detailed studies of the behavior of these structures, and how microgravity affects their unique ability to act like both a liquid and a solid crystal.

Treadmill 2 (T2) Yearly Maintenance: Lindgren performed a thorough inspection of T2's drive shaft, pulley teeth, motor, motor carriage bolts, motor belt, mounting screws, lateral belts, and clean the bearing housing. All observations were photo documented and downlinked to the ground for analysis. Once the maintenance was completed, T2 was powered up and is back in operation.

USOS Hatch Seal Inspections: Lindgren cleaned and inspected USOS Hatch Seals, Hatch Plate Sealing Surfaces, and Hatch Handle Mechanisms. Hatches targeted for today's inspection included the Node 1 (Port, Stbd, Fwd, Aft), Lab (Fwd, Aft), Node 2 (Aft), and Node 3 (Nadir, Fwd). Additional hatch inspections are scheduled to occur this Thursday, September 24th.

Sound Level Meter (SLM) Ops: Kelly measured the acoustic environment in habitable areas of the ISS by taking sound level readings. Today's measurements took place within Node 3, Cupola, MRM1, SM, and T2 operating at various speeds before the treadmill's yearly maintenance. One set of measurements, those following completion of the T2 yearly maintenance, were deferred to tomorrow due to crew time constraints.

Lab Bacteria Filter Change out: As part of routine maintenance, Yui changed out six bacteria filters in Lab today.

Water Recovery System (WRS) Total Organic Compound Analyzer (TOCA) Results: Today's TOC level was 2932 µg/L. This continues an upward trend, though currently still below the screening limit of 3000 µg/L. The WRS Ion Exchange (IX) Bed and Microbial Check Valve (MCV) replacement are planned in early October and should help lower the TOC levels. The date for the Multifiltration (MF) bed replacement is TBD.

Today's Planned Activities
All activities were completed unless otherwise noted.

EDV (KOV) (#1162 (00063575R, ФГБ1ПГО_4_404_1) Fill (Separation) from EDV #1165 (00063577R, ФГБ1ПГО_1_111). Label EDV "only for KOV") for Elektron or EDV-SV
Drive (start).
Ocular Health (ОН) - Vision Test
Ocular Health (ОН) - VIS Questionnaire
BIOME - Sample Collection
SM Ventilation System Cleaning (Group A)
Ocular Health (ОН) - Vision Test
СОЖ Maintenance
HRF - Sample MELFI Insertion
BIOME - Sample Stowage
OASIS - Sample Container Exchange
OH - Blood Pressure Measurement
Ocular Health (ОН) - VIS Questionnaire
Ocular Health (ОН) - Tonometry Hardware Setup
OH - Blood Pressure Measurement
BIOME - Sample Stowage
Ocular Health (ОН) - Tonometry Test
DAN. Experiment Ops.
DAN. Assistance during Experiment Ops
WRS - Water Samples Analysis
DAN. Photography during Experiment Ops
Ocular Health (ОН) - Tonometry Hardware Stow
Sound Level Meter (SLM) Data Transfer
ВД1 and ВД2 Air Ducts Vacuum Cleaning in DC1
DAN. Experiment Ops.
DAN. Assistance during Experiment Ops
Hatch Seal Inspection
DAN. Photography during Experiment Ops
Health Maintenance System (HMS) CBT
SEISMOPROGNOZ. Data Downlink from МКСД Hard Drive (end) and archiving
ISS Medical Accessories Kit (IMAK) Unpack from Soyuz No. 718 (44S)
КСПЭ Cables Inventory Audit
Post-Exercise Hygiene
Т2 - Yearly Maintenance
Video Recording of Greetings
Hardware Gather for LAB Bacteria Filter R&R
EXPRESS Rack 4 - Laptop Config
SM Ventilation System Cleaning (Group A)
MRM1 CTTC Configuration
VIZIR. Experiment Ops
EXPRESS Rack 4 - Laptop and Gateway Configuration
Removal of Hydraulic Connector from Replacement Unit No. 39 of the Hydraulic Circuit in MRM1.
LAB Bacteria Filter R&R
MELFI2 Kits Operations
[ВКС] Hardware Bag Search
LAB Bacteria Filter R&R
Ice Brick - Stow
Determine the contents of and relocate hardware from JPM1F6_E1.
Return to Nominal Comm Config
IMS Data File Prep
Hardware Restow after LAB Bacteria Filter R&R
TOCA - Data Recording
Sound Level Meter (SLM) Setup and Measurement Ops [Deferred]
PCS Relocate
Health Maintenance System (HMS) CBT
Sound Level Meter (SLM) Data Transfer [Deferred]
ПФ1 and ПФ2 Dust Filters R&R (ФГБ1ПГО_4_419_1, Bag 428-25 (00067328R). Update IMS) and [В1] and [В2] Fan Grills Cleaning in DC1
EKON-M. Observation and Photography
OPTIMIS Viewer Familiarization
Preparation of reports for Roscosmos site

Completed Task List Items
None

Ground Activities
All activities were completed unless otherwise noted.
Nominal System Commanding

Three-Day Look Ahead:
Tuesday, 09/22: 44S Emergency OBT, Ocular Health, HTV Cargo Ops, Cognition Test
Wednesday, 09/23: Sprint, Ocular Health, EVA Tether Inspect, EVA Safer Swap, RMCT Swap
Thursday, 09/24: Cardio Ox Ultrasound, Cell Biology Experiment, HTV Cargo Ops, Hatch Seal Inspect, RMCT Swap

QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Shutdown
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.